Apply the distributive property to the following expression, and then simplify the result:
20(-1/5x+1/20y)

Answer by ewatrrr(24785)


Hi
Apply the distributive property to the following expression, and then simplify
20(-1/5x+1/20y) = -4/x + 1/y = (-4y + x)/xy = (x-4y)/xy